# Topological regularity of Busemann spaces of nonpositive curvature

### v1

@@ 1. Introduction
The purpose of this paper is to extend recent topological results of Lytchak–Nagano [64], [65] and Lytchak–Nagano–Stadler [66] (building on earlier work of Thurston [90]) for CAT(00) spaces to the setting of BNPC spaces. This gives a positive answer to the problem raised in [66, Section 1.6]. As mentioned above, the difference between Alexandrov and Busemann curvature bounds reflects the difference between Riemannian and Finsler geometries. Therefore, the key to the extension is to capture the Finsler nature of BNPC spaces and deduce from it the same topological properties as in the CAT(00) case. Our technical contribution will be discussed in Section 1.6.
@@ 1.1. Main results
Theorems 1.1 and 1.2 generalize, respectively, [65, Theorems 1.1, 1.2] of Lytchak–Nagano in the CAT(00) case, and Theorem 1.3 generalizes [66, Theorem 1.1] of Lytchak–Nagano–Stadler. In Theorem 1.1, the homotopy type of a punctured ball is independent of its radius (and of its openness or closedness), thanks to the unique geodesic contraction centered at $p$. Although Theorem 1.1 may look different from [65, Theorem 1.1], they are essentially the same, because in the CAT(00) case a small punctured ball around $p$ is homotopy equivalent to the space of directions at $p$ ([58, Theorem A]). Note that Theorems 1.1 and 1.2 are local results, whereas Theorem 1.3 is a global one. For more generalizations, such as general curvature bounds, see Section 1.5.
@@ 1.2. Comments
In the situation of Theorem 1.1, the punctured ball around $p$ is homotopy equivalent to any sufficiently small metric sphere around $p$; see Theorem 1.5. However, for dimensions $\geq 5$, such a metric sphere need not be homeomorphic to a sphere, as a consequence of the double suspension theorem of Cannon [25] and Edwards [35] (cf. [14]). On the other hand, for dimension $\leq 4$, every sufficently small metric sphere around $p$ is a topological sphere, exactly as in the CAT(00) case (Corollary 7.3).
Theorem 1.2 for Alexandrov curvature bounds was a question raised by Quinn [86, Problem 7.2]. Note that for dimension $\geq 4$, a locally BNPC homology manifold is not necessarily a topological manifold. On the other hand, for dimension $\leq 3$, any locally BNPC homology manifold is a topological manifold. The dimension $1$ and $2$ cases are general facts holding without the BNPC condition (Theorem 2.10). The dimension $3$ case was essentially observed by Thurston [90, Theorem 3.3] (see Theorem 3.8). Our argument provides an alternative proof of this fact in Corollary 7.2.
@@ 1.3. General results
Similarly, in this paper we develop the theory of GNPC spaces. A GNPC space is a separable, locally compact, locally geodesically complete, locally BNPC space (the separability is not necessary for our purposes). Every GCBA space of nonpositive curvature is, in fact, a GNPC space. Throughout this paper, all spaces under consideration are GNPC spaces. We prove the following two theorems in this general setting; compare with [64, Theorems 1.2, 1.12] (cf. [56], [58]).
@@ 1.4. Applications
Andreev [7] later proved a more rigid result: the tangent cone of a locally BNPC G-space is a strictly convex normed space. In other words, any GNPC space without branching geodesics admits a Finsler-like structure. This can be viewed as a model case of the geometric structure of the “regular part” of a GNPC space. Compare with the Riemannian structure of the regular part of a GCBA space: [64, Theorem 1.3], [15]. See Section 8 for more details.
As an application of Theorem 1.1, we obtain the following topological stability theorem for locally BNPC G-spaces. This partially generalizes [65, Theorem 1.3]. The injectivity radius of a G-space $X$ is the supremum of $r>0$ such that for any $p\in X$ every shortest path emanating from $p$ of length $<r$ extends uniquely to length $r$ (i.e., without branching).
@@ 1.5. Generalizations
The other is a potential generalization to general curvature bounds. The original statements of Lytchak–Nagano and Lytchak–Nagano–Stadler were established for CAT($\kappa$) spaces with arbitrary $\kappa\in\mathbb{R}$ (even a global result; see [66, Corollary 1.8]). By using triangle comparison, Busemann curvature bounds can also be defined for any $\kappa\in\mathbb{R}$ (Remark 2.5). While it should be possible to extend our arguments to that broader setting, we do not pursue this here and leave it as an open problem (Problem 8.1).
@@ 1.6. Main tool
Acknowledgments. The authors are especially grateful to Stephan Stadler for giving us the opportunity to collaborate and for his insightful suggestions. The first author would like to thank Shin-ichi Ohta and Koichi Nagano for helpful discussions. The second author would like to thank Wenyuan Yang for his inquiry regarding potential applications to spaces admitting convex geodesic bicombings, and Shmuel Weinberger for helpful discussions on homology manifolds. The first author was supported by JSPS KAKENHI Grant Number 22KJ2099. The second author was supported by NSFC grant 12201102.
@@ 2.3. BNPC spaces
Indeed, Condition (2.1) is equivalent to the definition of a BNPC space ([79, Proposition 8.1.2]). The above two properties are, in a sense, closely related to the two angles we will introduce in Sections 4.1 and 4.2, respectively.
@@
  • Strictly convex Banach spaces ([79, Proposition 8.1.6]).
  • Complete, simply connected Finsler manifolds with reversible Berwald metrics of nonpositive flag curvature ([59], [60], [51], [74, Section 8.4]).
@@
  • CAT(00) cube complexes equipped with piecewise $L^{p}$-metrics, where $1<p<\infty$ ([48, Theorem D])
Note that if $(X,d)$ is BNPC, then its rescaling $(X,\lambda d)$ remains BNPC for any $\lambda>0$. The following are basic properties of BNPC spaces that immediately follow from the above definition (see [79, Chapter 8]).
@@
The above difference between BNPC spaces and CAT(00) spaces in a sense resembles the difference between CD spaces and RCD spaces, which are synthetic generalizations of manifolds with lower Ricci curvature bounds. Although the CD condition characterizes lower Ricci curvature bounds for Riemannian manifolds, it also makes sense for Finsler manifolds. To exclude Finsler-type spaces, one needs the so-called infinitesimally Hilbertian condition, which led to the notion of RCD space. See [74, Chapter 18] and the references therein.
@@ 2.7. Hurewicz fibrations
We recall the following condition from [65, Definition 4.3].
@@
We say that a map $f:X\to Y$ between topological spaces has a locally uniformly contractible fibers if the following holds: for any $x\in X$ and every neighborhood $U$ of $x$, there exists a neighborhood $V\subset U$ of $x$ such that for any fiber $\Pi$ of $f$ intersecting $V$, the intersection $\Pi\cap V$ is contractible in $\Pi\cap U$.
The following theorems of Ungar [94, Theorems 1, 2] provide sufficient conditions for a map to be a Hurewicz fibration (see also [65, Theorems 4.5, 4.6]).
@@
The following theorem of Raymond [87, Theorem (1)] describes the relationship between homology manifolds and Hurewicz fibrations (see also [65, Theorem 4.10]).
@@ 2.8. Recognition theorems
In the proofs of the main theorems, the key is to “recognize” topological manifolds among homology manifolds (cf. Theorem 2.10) and fiber bundles among Hurewicz fibrations. Our proofs rely heavily on such recognition theorems listed in [65, Sections 4.3, 6.1]. However, these recognition theorems are used in the topological part of the proofs, match exactly the CAT(00) case and thus omitted here. For this reason, although the use of these theorems is essential, we avoid including their statements and numerous related references here. See [65, Sections 4.3, 6.1] and the references therein.
@@ 3. Facts
The first proposition is about the finite-dimensionality of GNPC spaces ([64, Proposition 5.1, Corollary 5.4]). A metric space $X$ is called doubling, or more specifically $L$-doubling, if there exists a doubling constant $L>0$ such that every $r$-ball in $X$ can be covered by at most $L$ balls of radius $r/2$. Clearly, if any $r$-ball in $X$ contains at most $L$ points with pairwise distance $\geq r/2$, then $X$ is $L$-doubling (we say that such points with pairwise distance $\geq r/2$ are $r/2$-separated). A metric space is locally doubling if each point has a neighborhood that is doubling, where the doubling constant may depend on each point.
@@
The proof of the first statement is exactly the same as [64, Proposition 5.1]. Let $B$ be a tiny ball centered at $p\in X$. By rescaling, we may assume $B=B(p,1)$ (note that the doubling constant is a scaling invariant).
@@
It is known that if in addition $X$ is (globally) BNPC and doubling, then it admits a bi-Lipschitz embedding into finite-dimensional Euclidean space: see [62] and [98]. Compare with the GCBA case [64, Proposition 5.3].
@@
For GCBA spaces, the Hausdorff dimension is equal to the topological dimension, both locally and globally ([64, Theorem 1.1]). For now the authors do not know whether the same holds true for GNPC spaces (Problem 8.2). Note that this claim is false without geodesic completeness. A counterexample is a tree branching very fast, described in [56, P. 411, Footnote 3]. See also [56, Theorem D] for a related result under cocompact group actions.
The second proposition provides a sufficient condition for a BNPC space to be geodesically complete ([65, Lemma 2.2], [91, Proposition 2.1]). In particular, any (locally) BNPC space in the main theorems are (locally) geodesically complete.
@@
The third proposition addresses the local topology of BNPC homology manifolds ([65, Lemma 3.2], [90, Proposition 2.7]).
@@
The fourth theorem is the $3$-dimensional version of Theorems 1.2 and 1.3. The first half is due to Thurston [90, Theorem 3.3] and its second half follows from Rolfsen [88, Theorem 1] and Brown [19, Theorem]. Later we will provide an alternative proof of the first half (Corollary 7.2).
@@
The second half is an immediate consequence of the results of Rolfsen [88, Theorem 1] and Brown [19, Theorem]. Note that our BNPC metric is SC in the sense of Rolfsen. (The resolution of the Poincaré conjecture, together with the generalized Schoenflies theorem, also implies that any compact contractible $3$-manifold with boundary is a $3$-cell.) ∎
@@
In dimension $\geq 3$, many contractible open manifolds are not homeomorphic to Euclidean space. See for instance [46]. See also [66, Sections 1.2, 1.3] and references therein for geometric structures of such spaces.
@@ 4.1. Almost nonbranching
This angle controls the branching of geodesics emanating from $p$. In the GCBA case, the independence of this angle on parametrizations implies that the tangent cone (i.e., the Gromov–Hausdorff blow-up) at $p$ is a Euclidean cone ([64, Corollary 5.7]). Since the geodesics from the vertex of the Euclidean cone are nonbranching, one sees that the geodesics from $p$ are almost nonbranching near $p$ ([64, Proposition 7.3]). In the GNPC case, we must restrict this argument to some almost fixed scale. The following is a uniform version of the convergence in Definition 4.3 with respect to some almost fixed scale.
@@
This follows from local compactness and the Busemann monotonicity (2.1). The proof is almost the same as in the GCBA case, [64, Lemma 5.6], except that we have to restrict our argument to some almost fixed scale as in the assumption. In what follows, we denote by $\varkappa(\delta)$ various positive functions such that $\varkappa(\delta)\to 0$ as $\delta\to 0$, as explained in Notation 4.1.
@@ 4.2. Almost orthogonality
Indeed, it might be better to consider the differential quotient $t^{-1}(|px(t)|-|pq|)$ rather than the comparison angle ~$\tilde{\angle}pqx(t)$, since the former is really monotone whereas the latter is almost monotone. Nevertheless, we stick to the latter angle notation for inspiring geometric intuition and for comparing with the GCBA case.
@@
Combining the above three lemmas, we obtain the following key property. In the GCBA case, this is a combination of the triangle inequality and the upper semicontinuity of angles ([64, Lemma 5.5]).
@@ 5.1. Definition and existence
In this subsection, we define strainers for GNPC spaces and show their existence (cf. [64, Section 7]).
@@
viewed from $p$ at $q$ toward $x$, introduced in Section 4.2. Recall that this angle was essentially the directional derivative of the distance function from $p$, i.e.,
@@
where $x(t)$ is the unit-speed shortest path from $q$ to $x$ (see Remark 4.10). We note that the other angle introduced in Section 4.1 is not used for the definition of a strainer, but will be used for the existence later.
@@
Now we prove the existence of a strainer. First we show the existence of a $1$-strainer. This follows from the key lemmas in Section 4.1, specifically Proposition 4.6. Compare with [64, Proposition 7.3].
@@
Next we show the existence of a $k$-strainer, or more precisely, how to extend a strainer. This follows from the previous proposition and the key lemmas in Section 4.2, specifically Proposition 4.14. Compare with [64, Proposition 9.4].

### v2

@@ 1. Introduction
The purpose of this paper is to extend recent topological results of Lytchak–Nagano [63], [64] and Lytchak–Nagano–Stadler [65] (building on earlier work of Thurston [90]) for CAT(00) spaces to the setting of BNPC spaces. This gives a positive answer to the problem raised in [65, Section 1.6]. As mentioned above, the difference between Alexandrov and Busemann curvature bounds reflects the difference between Riemannian and Finsler geometries. Therefore, the key to the extension is to capture the Finsler nature of BNPC spaces and deduce from it the same topological properties as in the CAT(00) case. Our technical contribution will be discussed in Section 1.6.
@@ 1.1. Main results
Theorems 1.1 and 1.2 generalize, respectively, [64, Theorems 1.1, 1.2] of Lytchak–Nagano in the CAT(00) case, and Theorem 1.3 generalizes [65, Theorem 1.1] of Lytchak–Nagano–Stadler. In Theorem 1.1, the homotopy type of a punctured ball is independent of its radius (and of its openness or closedness), thanks to the unique geodesic contraction centered at $p$. Although Theorem 1.1 may look different from [64, Theorem 1.1], they are essentially the same, because in the CAT(00) case a small punctured ball around $p$ is homotopy equivalent to the space of directions at $p$ ([56, Theorem A]). Note that Theorems 1.1 and 1.2 are local results, whereas Theorem 1.3 is a global one. For more generalizations, such as general curvature bounds, see Section 1.5.
@@ 1.2. Comments
In the situation of Theorem 1.1, the punctured ball around $p$ is homotopy equivalent to any sufficiently small metric sphere around $p$; see Theorem 1.5. However, for dimensions $\geq 5$, such a metric sphere need not be homeomorphic to a sphere, as a consequence of the double suspension theorem of Cannon [25] and Edwards [35] (cf. [14]). On the other hand, for dimension $\leq 4$, every sufficiently small metric sphere around $p$ is a topological sphere, exactly as in the CAT(00) case (Corollary 7.3).
Theorem 1.2 for Alexandrov curvature bounds was a question raised by Quinn [85, Problem 7.2]. Note that for dimension $\geq 4$, a locally BNPC homology manifold is not necessarily a topological manifold. On the other hand, for dimension $\leq 3$, any locally BNPC homology manifold is a topological manifold. The dimension $1$ and $2$ cases are general facts holding without the BNPC condition (Theorem 2.10). The dimension $3$ case was essentially observed by Thurston [90, Theorem 3.3] (see Theorem 3.9). Our argument provides an alternative proof of this fact in Corollary 7.2.
@@ 1.3. General results
Similarly, in this paper we develop the theory of GNPC spaces. A GNPC space is a separable, locally compact, locally geodesically complete, locally BNPC space (the separability is not necessary for our purposes). Every GCBA space of nonpositive curvature is, in fact, a GNPC space. Throughout this paper, all spaces under consideration are GNPC spaces. We prove the following two theorems in this general setting; compare with [63, Theorems 1.2, 1.12] (cf. [54], [56]).
@@ 1.4. Applications
Andreev [7] later proved a more rigid result: the tangent cone of a locally BNPC G-space is a strictly convex normed space. In other words, any GNPC space without branching geodesics admits a Finsler-like structure. This can be viewed as a model case of the geometric structure of the “regular part” of a GNPC space. Compare with the Riemannian structure of the regular part of a GCBA space: [63, Theorem 1.3], [15]. See Section 8 for more details.
As an application of Theorem 1.1, we obtain the following topological stability theorem for locally BNPC G-spaces. This partially generalizes [64, Theorem 1.3]. The injectivity radius of a G-space $X$ is the supremum of $r>0$ such that for any $p\in X$ every shortest path emanating from $p$ of length $<r$ extends uniquely to length $r$ (i.e., without branching).
@@ 1.5. Generalizations
The other is a potential generalization to general curvature bounds. The original statements of Lytchak–Nagano and Lytchak–Nagano–Stadler were established for CAT($\kappa$) spaces with arbitrary $\kappa\in\mathbb{R}$ (even a global result; see [65, Corollary 1.8]). By using triangle comparison, Busemann curvature bounds can also be defined for any $\kappa\in\mathbb{R}$ (Remark 2.5). While it should be possible to extend our arguments to that broader setting, we do not pursue this here and leave it as an open problem (Problem 8.1).
@@ 1.6. Main tool
Acknowledgments. The authors are especially grateful to Stephan Stadler for giving us the opportunity to collaborate and for his insightful suggestions. The first author would like to thank Shin-ichi Ohta and Koichi Nagano for helpful discussions, and Alexander Lytchak for his valuable comments, especially on Section 8. The second author would like to thank Wenyuan Yang for his inquiry regarding potential applications to spaces admitting convex geodesic bicombings, and Shmuel Weinberger for helpful discussions on homology manifolds. The first author was supported by JSPS KAKENHI Grant Number 22KJ2099. The second author was supported by NSFC grant 12201102.
@@ 2.3. BNPC spaces
Indeed, Condition (2.1) is equivalent to the definition of a BNPC space, thanks to the triangle inequality ([78, Proposition 8.1.2]). The above two properties are, in a sense, closely related to the two angles we will introduce in Sections 4.1 and 4.2, respectively.
@@
  • Strictly convex Banach spaces ([78, Proposition 8.1.6]).
  • Complete, simply connected Finsler manifolds with reversible Berwald metrics of nonpositive flag curvature ([57], [58], [51], [73, Section 8.4]).
@@
  • CAT(00) cube complexes equipped with piecewise $L^{p}$-metrics, where $1<p<\infty$ ([48, Theorem D]).
Note that if $(X,d)$ is BNPC, then its rescaling $(X,\lambda d)$ remains BNPC for any $\lambda>0$. The following are basic properties of BNPC spaces that immediately follow from the above definition (see [78, Chapter 8]).
@@
The above difference between BNPC spaces and CAT(00) spaces in a sense resembles the difference between CD spaces and RCD spaces, which are synthetic generalizations of manifolds with lower Ricci curvature bounds. Although the CD condition characterizes lower Ricci curvature bounds for Riemannian manifolds, it also makes sense for Finsler manifolds. To exclude Finsler-type spaces, one needs the so-called infinitesimally Hilbertian condition, which led to the notion of RCD space. See [73, Chapter 18] and the references therein.
@@ 2.7. Hurewicz fibrations
We recall the following condition from [64, Definition 4.3].
@@
We say that a map $f:X\to Y$ between topological spaces has locally uniformly contractible fibers if the following holds: for any $x\in X$ and every neighborhood $U$ of $x$, there exists a neighborhood $V\subset U$ of $x$ such that for any fiber $\Pi$ of $f$ intersecting $V$, the intersection $\Pi\cap V$ is contractible in $\Pi\cap U$.
The following theorems of Ungar [94, Theorems 1, 2] provide sufficient conditions for a map to be a Hurewicz fibration (see also [64, Theorems 4.5, 4.6]).
@@
The following theorem of Raymond [86, Theorem (1)] describes the relationship between homology manifolds and Hurewicz fibrations (see also [64, Theorem 4.10]).
@@ 2.8. Recognition theorems
In the proofs of the main theorems, the key is to “recognize” topological manifolds among homology manifolds (cf. Theorem 2.10) and fiber bundles among Hurewicz fibrations. Our proofs rely heavily on such recognition theorems listed in [64, Sections 4.3, 6.1]. However, these recognition theorems are used in the topological part of the proofs, match exactly the CAT(00) case and thus omitted here. For this reason, although the use of these theorems is essential, we avoid including their statements and numerous related references here. See [64, Sections 4.3, 6.1] and the references therein.
@@ 3. Facts
The first proposition is about the finite-dimensionality of GNPC spaces ([63, Proposition 5.1, Corollary 5.4]). A metric space $X$ is called doubling, or more specifically $L$-doubling, if there exists a doubling constant $L>0$ such that every $r$-ball in $X$ can be covered by at most $L$ balls of radius $r/2$. Clearly, if any $r$-ball in $X$ contains at most $L$ points with pairwise distance $\geq r/2$, then $X$ is $L$-doubling (we say that such points with pairwise distance $\geq r/2$ are $r/2$-separated). A metric space is locally doubling if each point has a neighborhood that is doubling, where the doubling constant may depend on each point.
@@
The proof of the first statement is exactly the same as [63, Proposition 5.1]. Let $B$ be a tiny ball centered at $p\in X$. By rescaling, we may assume $B=B(p,1)$ (note that the doubling constant is a scaling invariant).
@@
The fact that $X$ is an ANR holds even without assuming local compactness or geodesic completeness; that is, any locally BNPC space is an ANR. Indeed, the proof of [56, Theorem 3.2] for locally CAT($\kappa$) spaces also applies to locally BNPC spaces. See also the references listed in [56, P. 353, Footnote 4].
Remark 3.3.
@@
It is known that if in addition $X$ is (globally) BNPC and doubling, then it admits a bi-Lipschitz embedding into finite-dimensional Euclidean space: see [60] and [98]. Compare with the GCBA case [63, Proposition 5.3].
@@
For GCBA spaces, the Hausdorff dimension is equal to the topological dimension, both locally and globally ([63, Theorem 1.1]). For now the authors do not know whether the same holds true for GNPC spaces (Problem 8.2). Note that this claim is false without geodesic completeness. A counterexample is a tree branching very fast, described in [54, P. 411, Footnote 3]. See also [54, Theorem D] for a related result under cocompact group actions.
The second proposition provides a sufficient condition for a BNPC space to be geodesically complete ([64, Lemma 2.2], [91, Proposition 2.1]). In particular, any (locally) BNPC space in the main theorems are (locally) geodesically complete.
@@
The third proposition addresses the local topology of BNPC homology manifolds ([64, Lemma 3.2], [90, Proposition 2.7]).
@@
The fourth theorem is the $3$-dimensional version of Theorems 1.2 and 1.3. The first half is due to Thurston [90, Theorem 3.3] and its second half follows from Rolfsen [87, Theorem 1] and Brown [19, Theorem]. Later we will provide an alternative proof of the first half (Corollary 7.2).
@@
The second half is an immediate consequence of the results of Rolfsen [87, Theorem 1] and Brown [19, Theorem]. Note that our BNPC metric is SC in the sense of Rolfsen. (The resolution of the Poincaré conjecture, together with the generalized Schoenflies theorem, also implies that any compact contractible $3$-manifold with boundary is a $3$-cell.) ∎
@@
In dimension $\geq 3$, many contractible open manifolds are not homeomorphic to Euclidean space. See for instance [46]. See also [65, Sections 1.2, 1.3] and references therein for geometric structures of such spaces.
@@ 4.1. Almost nonbranching
This angle controls the branching of geodesics emanating from $p$. In the GCBA case, the independence of this angle on parametrizations implies that the tangent cone (i.e., the Gromov–Hausdorff blow-up) at $p$ is a Euclidean cone ([63, Corollary 5.7]). Since the geodesics from the vertex of the Euclidean cone are nonbranching, one sees that the geodesics from $p$ are almost nonbranching near $p$ ([63, Proposition 7.3]). In the GNPC case, we must restrict this argument to some almost fixed scale. The following is a uniform version of the convergence in Definition 4.3 with respect to some almost fixed scale.
@@
This follows from local compactness and the Busemann monotonicity (2.1). The proof is almost the same as in the GCBA case, [63, Lemma 5.6], except that we have to restrict our argument to some almost fixed scale as in the assumption. In what follows, we denote by $\varkappa(\delta)$ various positive functions such that $\varkappa(\delta)\to 0$ as $\delta\to 0$, as explained in Notation 4.1.
@@ 4.2. Almost orthogonality
Indeed, it might be better to consider the differential quotient $t^{-1}(|px(t)|-|pq|)$ rather than the comparison angle ~$\tilde{\angle}pqx(t)$, since the former is really monotone whereas the latter is almost monotone. Although we stick to the latter angle notation for inspiring geometric intuition and for comparing with the GCBA case, the above viewpoint plays an important role in the definition of a strainer in the next section.
@@
Combining the above three lemmas, we obtain the following key property. In the GCBA case, this is a combination of the triangle inequality and the upper semicontinuity of angles ([63, Lemma 5.5]).
@@ 5.1. Definition and existence
In this subsection, we define strainers for GNPC spaces and show their existence (cf. [63, Section 7]).
@@
viewed from $p$ at $q$ toward $x$, introduced in Section 4.2. We again emphasize that this angle is essentially the directional derivative of the distance function from $p$, i.e.,
@@
where $x(t)$ is the unit-speed shortest path from $q$ to $x$ (see Remark 4.10). Therefore, the reader should consider the following definition of a strainer as controlling the derivative of the distance function rather than the angles, unlike in the GCBA case. Note that the other angle introduced in Section 4.1 is not used for the definition of a strainer, but will be used later to show the existence of a strainer.
@@
Now we prove the existence of a strainer. First we show the existence of a $1$-strainer. This follows from the key lemmas in Section 4.1, specifically Proposition 4.6. Compare with [63, Proposition 7.3].
@@
Next we show the existence of a $k$-strainer, or more precisely, how to extend a strainer. This follows from the previous proposition and the key lemmas in Section 4.2, specifically Proposition 4.14. Compare with [63, Proposition 9.4].